10 people can build 2 houses in 4 days.
how many people are required to build 5 houses in 2 days?
show process and decison making.
Answer #1

Ans:

Given that 10 people can build 2 houses in 4 days,then number of people required to build 5 houses in 2 days.

10*2*4=n*5*2

80=n*10

n=8

8 number of people are required to build 5 houses in 2 days.